Problem p2 SNF 2009-03-10 16:29:22: sensor head does not move

gsosa at snf.stanford.edu gsosa at snf.stanford.edu
Fri Mar 27 14:54:27 PDT 2009


System came back from repair. Re-installed and tested system-All OK. Vendors repaired the sensor head and did a PM. Uli tested system with calibration wafer- OK. Please note that the power is not to be turned off anymore. See note near the power switch that states : New  Leave power on 3/27/09. Also follow the new instructions that will be placed at the tool. If you have any questions, contact Uli System is ready to use.




More information about the p2-pcs mailing list